OpenStudy (anonymous):

Use a calculator to find the value of the expression. Then show that the calculated value is correct. frac{sqrt{2}+sqrt{6}}{sqrt{2+sqrt{3}}} =2 ~~I have tried subbing in a common variable to prove this, such as m= sqrt{2} and using that, put it into the equation so that the equation now reads: frac{m+(m^2+4)^{1/2}}{sqrt{2+(m^2+1)^{1/2}}} I have not gotten very far with that though, I may be doing it completely wrong.So any insight/help is appreciated! Thanks!

OpenStudy (anonymous):

\[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{2+\sqrt{3}}} =2 \] is the original equation.

OpenStudy (anonymous):

The equation I have come up with (when I sub in m as equaling\[\sqrt{2}\] is: \[\frac{m+(m^2+4)^{1/2}}{\sqrt{2+(m^2+1)^{1/2}}} \]

OpenStudy (anonymous):

the values comes around 4.8 use your calculator again

OpenStudy (anonymous):

...Oh wait you mean the calculated answer of the original equation isn't 2?..let me check..

OpenStudy (anonymous):

hmm... for this equation \[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{2+\sqrt{3}}} \] is still got it equaling 2..

OpenStudy (anonymous):

oh sorry i got denominator (2)^1/2 + 3^1/2 ....i'm on your question thanks :)

OpenStudy (anonymous):

Oh not a problem! :)

OpenStudy (anonymous):

Hey I got it !!!

OpenStudy (anonymous):

No way!! I spent like 2 hours on this today!

OpenStudy (anonymous):

2^1/2 +6^1/2 y = ------------- ( 2 + 3^1/2)^1/2 now just square both sides

OpenStudy (anonymous):

2 + 6 + 2. (12)^1/2 y^2 = ----------------- 2 + 3^1/2 8 + 2 . 12^1/2 y^2 = ------------- 2 + 3^1/2

OpenStudy (anonymous):

2( 4 + (4.3)^1/2) y ^2 = ------------ 2 + 3^1/2

OpenStudy (anonymous):

2 * 4^1/2 ( 4^1/2 + 3^1/2) y^2 = ---------------------------- 2 + 3^1/2 y^2 = 4.( 2 + 3^1/2) -------------- 2 + 3^1/2 y^2 = 4 y = 2

OpenStudy (anonymous):

I squared just because it helps get rid of that whole under root .... and it makes easy to think ...

OpenStudy (anonymous):

Okay! I getcha~ my only question is after you square both sides where do you get the 2. (12)^1/2 from?

OpenStudy (anonymous):

(12)^1/2 = (4.3)^1/2 = (4)^1/2 . (3)^1/2 = 2 . (30^1/2

OpenStudy (anonymous):

Oh okay! Well thank you so much Ishaan!! That makes sense! ( I think I was just overthinking the whole thing! ) :)

OpenStudy (anonymous):

: )